# Two-Player Zero-Sum Games with Bandit Feedback

### v3

@@ Abstract
We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback. We propose three algorithms based on the Explore-Then-Commit (ETC) framework. The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration. Our objective is to demonstrate the applicability of ETC in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that ETC-based algorithms perform effectively in zero-sum game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
Keywords zero-sum games, regret analysis, bandit feedback
@@ 1 Introduction
This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in [33, 34]. These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning. In such settings, players repeatedly select actions from finite sets, with payoffs determined by a fixed but unknown game matrix.
Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix [7]. In this paper, we study a TPZSG setting with bandit feedback, where players observe each other’s actions, by learning pure strategies for both players through balancing of exploration and exploitation.
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for adversarial games under bandit feedback has been analyzed using worst-case regret bounds in [29]. In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated in TPZSGs [18, 24]. However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs. While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to zero-sum game settings, which makes this the first work to provide its detailed analysis in such a setting. A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure. Moreover, our main focus is to show whether ETC and action elimination algorithms is applicable to zero-sum games by offering a refined understanding of regret behavior in TPZSGs under bandit feedback.
In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses. The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ . The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players. Finally, the last algorithm extends the adaptive elimination approach through non-uniform exploration, which also achieves a regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Consequently, despite the simplicity of the ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
Paper structure.
Section 2 reviews related work. Section 3 introduces preliminaries and formalizes a two-player zero-sum game setting. In Section 4, we present the ETC in a TPZSG setting and analyze its regret. Section 5 introduces an adaptive elimination algorithm and provides its regret analysis. In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret upper bound. Section 7 offers empirical results to validate their theoretical performances. Finally, Section 8 discusses our conclusions and future research directions.
@@ 2 Related Work
The multi-armed bandit (MAB) problem, has been introduced by [32] and originally described by [31], has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as [3, 4, 5, 6, 8, 9]. The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has been formalized by [19] and further explored in studies such as [2, 21].
@@
Zero-sum games represent a core concept in game theory. They describe competitive scenarios in which the gain of one player is exactly equal to the loss of the other. First highlighted by [33, 34], their theory and applications have been extensively explored [8, 12, 35]. The Minimax Theorem in [33, 34] guarantees the existence of equilibrium in TPZSGs, but finding equilibrium becomes significantly harder when the payoff matrix is unknown [13, 14, 25, 37].
@@
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations. Various approaches, such as UCB [29], Tsallis-INF [18] and a novel algorithm [24], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments. Additionally, [23] and [22] investigate instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
While [29] focuses on the worst-case regret analysis and [24] analyzes instance-dependent regret only for a $2\times 2$ game matrix, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium. Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action. Since their regret bound involves the square of the sum of suboptimality gaps, our instance-dependent analysis gives a tighter bound.
@@ 3 Preliminaries
In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on, and the gaps $\Delta$ that we use for the analysis.
In a TPZSG, the goals of the players are strictly opposed, which implies that any gain by one player corresponds to an equal loss by the other. This framework provides a fundamental model for adversarial interactions where strategic decision-making under competition is central. The outcome of such a game is fully determined by the strategies chosen by both players and the goal of each player is to maximize their own payoff while minimizing that of their opponent.
@@ 3.1 Two-Player Zero-Sum Games
$S_{x}$ and $S_{y}$ are the action sets for the row player and column player, respectively. Thus, the action space of the game is $S_{A}=S_{x}\times S_{y}$, so that the action pairs $(i,j)\in S_{A}$ where $i\in S_{x},j\in S_{y}$. We also define $m=|S_{x}|$ as the number of actions for the row player and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs. The game itself is defined through a game matrix $A$, therefore if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player A(i,j)-A(i,j).
The row player has a strategy (i.e. a probability distribution) $p$ to select an action $i\in S_{x}$, while the column player selects an action $j\in S_{y}$ according to a strategy $q$.
@@
$V^{*}$ $\geq{p}^{\top}Aq^{*},$ $V^{*}$ $\leq{p^{*}}^{\top}Aq$ (1)
where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game [26]. This means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
@@
A pair $(i^{*},j^{*})$ is a unique pure NE if:
$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$ (2)
or equivalently
$A(i,j^{*})\leq A(i^{*},j^{*})\leq A(i^{*},j),\qquad\forall i\in S_{x},j\in S_{y}.$ (3)
The condition in (3) ensures that no player has an incentive to deviate to another action. We use $V^{*}=A(i^{*},j^{*})$ to denote the value of the game.
@@ 3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$. The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, where the maximizing player receives $r_{t}$ and the minimizing player receives $r_{t}$. We assume that both players observe each other’s action.
Since the players do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. We express the action pair played in round $t$ by $(i_{t},j_{t})$. In each round $t$, the estimated game matrix is updated by computing the average payoff for each action pair $(i,j)$ as follows:
@@ 3.3 Gap Notions
Using the NE property in (3), for all $i\in S_{x}$ and $j\in S_{y}$ we can write the following:
$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)$
which implies
@@
Moreover,
$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$
implies the following:
@@
Then, the inequality (9) implies that $\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$ and similarly from (10), we have $\Delta_{ij}^{*}\geq-\Delta_{ij}^{\text{min}}$. Therefore, we have $\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$. On the other hand, since $\Delta_{ij}^{\text{min}}\geq 0$, we can write
$\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}+\Delta_{ij}^{\text{min}}=\Delta_{ij}.$ (11)
@@ 3.4 Regret Notions
$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i,j_{t})-A(i_{t},j_{t})\Big]$
$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\Big]$
@@
The (combined) external regret is given by the following:
@@
External regret quantifies how much worse the players have performed compared to their best actions in hindsight. An alternative regret notion is Nash regret, which measures the deviation of value from the Nash equilibrium:
@@ 4 ETC-TPZSG Algorithm
Applying ETC in a zero-sum game setting presents some challenges compared to the standard MAB scenario. In bandit problems, after the exploration phase, a player simply commits to the arm with the highest estimated reward, which is typically effective if exploration is sufficient. However, in a zero-sum game with pure NE, the objective is not to identify an action with high reward but to find a possibly true NE. Inadequate exploration can lead to poor estimates of payoffs, making it difficult to correctly identify such an equilibrium. As a result, the design of the ETC approach in the zero-sum game setting requires careful balance of the exploration phase.
@@
8:  Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
@@
Theorem 1.
The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs and $1\leq Nk\leq T$, is upper bounded as follows:
$R_{T}^{*}\leq k\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}+(T-Nk)\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}$ (14)
where $k$ is the exploration time per action pair and $N$ is the total number of action pairs.
Proof Sketch.
The total regret consists of two components: the regret from exploration and the regret from exploitation. During the exploration phase, each action pair is played $k$ times. In the exploitation phase, the algorithm plays according to the pure equilibria, where we account for the probability that the pure NE is not correctly identified. In other words, in the exploitation phase, we analyze the loss incurred from playing with a misidentified NE. Thus, the expected number of plays for any suboptimal action pair $(i,j)$ can be written by the following:
$\mathbb{E}[n_{ij,T}]=k+(T-Nk)\ \mathbb{P}((i,j)\text{ is identified as NE})$
Once this probability is bounded, the proof is complete. A detailed derivation is provided in Appendix B. ∎
@@
$R_{T}^{*}$ $\leq k\Delta+(T-2k)\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}$ (15)
$\leq k\Delta+T\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}.$ (16)
@@
$k=\max\Big\{1,\Big\lceil\frac{16\sigma^{2}}{\Delta^{2}}\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\rceil\Big\}$ (17)
@@
$R_{T}^{*}\leq\min\Big\{T\Delta,\Delta+\frac{16\sigma^{2}}{\Delta}\Big(1+\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}\Big)\Big\}$ (18)
where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (16) with the value of the exploration time in (17). Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we obtain $\Delta\geq\frac{4\sigma}{\sqrt{T}}$. Thus, when we put it in the regret bound (18), we have the following:
@@
where $c>0$ is some constant. If $\Delta\leq 1$, it becomes $R_{T}^{*}\leq 1+c\sqrt{T}$, which refers to an instance-independent bound since it depends only on the time horizon $T$ not on the game instance. In the zero-sum game setting, our analysis shows that we get a regret bound similar to the standard bandit setting in [20], indicating that the ETC-based algorithm performs effectively in adversarial environments.

### v4

@@ Abstract
We study a two-player zero-sum game in which the row player aims to maximize their payoff against a competing column player, under an unknown payoff matrix estimated through bandit feedback. We propose three algorithms based on the Explore-Then-Commit (ETC) and action pair elimination frameworks. The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration. Our objective is to demonstrate the applicability of ETC and action pair elimination algorithms in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O\left(\frac{\log(T\Delta^{2})}{\Delta}\right)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that the ETC and action pair elimination algorithms perform effectively in zero-sum game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
@@ 1 Introduction
This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in a setting with strictly opposing objectives, as first analyzed in [33, 34]. These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning. In such settings, players repeatedly select actions from finite sets, with payoffs determined by a fixed but unknown game matrix.
Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix [7]. Zero-sum games with bandit feedback setting, where players also observe each other’s actions, were introduced in [29]. In this paper, we study a TPZSG setting under bandit feedback with centralized control over both players by learning pure strategies through balancing of exploration and exploitation.
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for zero-sum games under bandit feedback has been analyzed using worst-case regret bounds in [29]. However, worst-case ignores the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated in TPZSGs [18, 24]. However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) and action elimination algorithms in TPZSGs. While ETC has been widely studied in standard multi-armed bandit problems, there has been no prior work specifically analyzing its performance to zero-sum game settings. Thus, our work is the first to provide a detailed analysis of ETC in this context. A key motivation for studying ETC is its algorithmic simplicity, which enables a clear structure. Our primary objective is to determine whether ETC and action elimination algorithms are applicable to zero-sum games by offering a refined understanding of regret behavior.
In particular, we propose three algorithms that adapt the ETC and action pair elimination approaches to the TPZSG setting and provide their instance-dependent regret analyses. The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieving a regret upper bound of $O(\Delta+\sqrt{T})$ (Section 4). The second algorithm incorporates adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of an $\varepsilon$-Nash Equilibrium; this yields an upper regret bound of $O\left(\frac{\log(T\Delta^{2})}{\Delta}\right)$ (Section 5). Intuitively, this approach is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players. Finally, the third algorithm extends the adaptive elimination approach through non-uniform exploration, also achieving the same regret bound (see Section 6). Consequently, our results demonstrate that ETC and elimination approaches can be highly effective in zero-sum game settings.
@@ 2 Related Work
The multi-armed bandit (MAB) problem was introduced by [32] and originally described by [31]. It has since become a crucial framework for online decision-making under uncertainty, as explored in works such as [3, 4, 5, 6, 8, 9]. The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff. This formalization was established by [19] and further investigated in studies such as [2, 21].
@@
Zero-sum games represent a core concept in game theory, which describe competitive scenarios in which the gain of one player is exactly equal to the loss of the other. First highlighted by [33, 34], their theory and applications have been extensively explored [8, 12, 35]. The Minimax Theorem in [33, 34] guarantees the existence of equilibrium in TPZSGs, but finding equilibrium becomes significantly harder when the payoff matrix is unknown [13, 14, 25, 37].
@@
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without knowledge of the true payoff matrix. Various approaches, such as UCB [29], Tsallis-INF [18] and a novel algorithm [24], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in such environments. In addition, [23] and [22] investigate instance-dependent sample complexity bounds while [11] examines the convergence rates to NE.
While [29] focuses on the worst-case regret analysis, achieving $O(\sqrt{T})$, and [24] analyzes instance-dependent regret only for a $2\times 2$ game matrix, [18] provides an instance-dependent analysis achieving an upper bound of $O\left(\left(\sum_{i\neq i^{*}}\frac{1}{\Delta_{i}}+\sum_{j\neq j^{*}}\frac{1}{\Delta^{\prime}_{j}}\right)\log T\right)$ in the case of pure equilibria where $\Delta$ and $\Delta^{\prime}$ denote some suboptimality gaps. Consequently, developing a comprehensive understanding of instance-dependent regret in TPZSGs under bandit feedback remains insufficiently addressed. In contrast to [18], where both players independently follow the Tsallis-INF algorithm, our setting is closely related to [29] and [24], but we assume control over both players. Our goal is to demonstrate the applicability of ETC and action elimination algorithms to zero-sum games and to achieve comparable regret bounds.
@@ 3 Preliminaries
In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on, and the gaps $\Delta$ that we use for the analysis. In a TPZSG, the goals of the players are strictly opposed, which implies that any gain by one player corresponds to an equal loss by the other. The outcome of such a game is fully determined by the strategies chosen by players and the goal of each player is to maximize their own payoff while minimizing that of their opponent.
@@ 3.1 Two-Player Zero-Sum Games
Let $S_{x}$ and $S_{y}$ denote the action sets of the row and column players, respectively, with the joint action space $S_{A}=S_{x}\times S_{y}$. We define $m=|S_{x}|$ and $l=|S_{y}|$ as the numbers of actions available to the row and column players, and let $N=ml=|S_{A}|$ be the total number of action pairs. The game is specified by a payoff matrix $A$; therefore, if $(i,j)$ is played, the row player receives a payoff of $A(i,j)$ and the column player receives A(i,j)-A(i,j). Moreover, the row player has a strategy (i.e. a probability distribution) $p$ to select an action $i\in S_{x}$, while the column player selects an action $j\in S_{y}$ according to a strategy $q$.
@@
${p}^{\top}Aq^{*}\leq{p^{*}}^{\top}Aq^{*}\leq{p^{*}}^{\top}Aq$ (1)
where ${p^{*}}^{\top}Aq^{*}$ is the value of the game [26].
@@
A pair $(i^{*},j^{*})$ is a unique pure NE if $i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)$ and $j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j)$ or equivalently
$A(i,j^{*})\leq A(i^{*},j^{*})\leq A(i^{*},j),\quad\forall i\in S_{x},j\in S_{y}.$ (2)
The condition in (2) ensures that no player has an incentive to deviate to another action.
@@ 3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from $q_{t}$. The players then observe a noisy payoff $r_{t}$ with $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, where the maximizing player receives $r_{t}$ and the minimizing player receives $r_{t}$. We assume that both players observe each other’s action. [29] introduced this setting with algorithms for row player against any column player; we instead assume centralized control over both players.
Since the players do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. In each round $t$, the estimated game matrix is updated by computing the average payoff for each action pair $(i,j)$ as follows:
@@ 3.3 Gap Notions
Lemma 1.
For all $i\in S_{x}$ and $j\in S_{y}$, $\Delta_{ij}^{*}\leq\Delta_{ij}$.
Proof.
Using the NE property in (2), for all $i\in S_{x}$ and $j\in S_{y}$ we can write $A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)$, which implies
@@
Similary, since $A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$ we have the following:
@@
Then, the inequality (8) implies that $\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$ and similarly from (9), we have $\Delta_{ij}^{*}\geq-\Delta_{ij}^{\text{min}}$. Therefore, we have $\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$. Since $\Delta_{ij}^{\text{min}}\geq 0$, it concludes the proof. ∎
@@ 3.4 Regret Notions
$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\left[\sum_{t=1}^{T}A(i,j_{t})-A(i_{t},j_{t})\right]$
$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\left[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\right]$
@@
The (combined) external regret, which quantifies how much worse the players have performed compared to their best actions in hindsight, is given by the following:
@@
An alternative regret notion is Nash regret, which measures the value of the deviation from the Nash equilibrium:
@@ 4 ETC-TPZSG Algorithm
Theorem 1.
For $1\leq Nk\leq T$, the Nash regret of Algorithm 1, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
$R_{T}^{*}\leq k\sum_{i,j}\Delta_{ij}^{*}+(T-Nk)\sum_{i,j}\Delta_{ij}^{*}\exp{\left(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\right)}$ (12)
where $k$ is the exploration time per action pair $(i,j)\in S_{A}$ and $N$ is the total number of action pairs.
Proof Sketch.
The total regret consists of two components: the regret from exploration and the regret from exploitation. During the exploration phase, each action pair is played $k$ times. In the exploitation phase, the algorithm plays according to the estimated pure equilibria. In other words, in the exploitation phase, we analyze the loss incurred from playing with a misidentified NE. Thus, the expected number of plays for any suboptimal action pair $(i,j)$ can be written by $\mathbb{E}[n_{ij,T}]=k+(T-Nk)\ \mathbb{P}((i,j)\text{ is identified as NE})$. Once this probability is bounded, the proof is complete. (See Appendix B). ∎
@@
8:  Play an equilibrium $(i^{*},j^{*})$ satisfying:
@@
$R_{T}^{*}\leq k\Delta+T\Delta\exp{\left(-\frac{k\Delta^{2}}{16\sigma^{2}}\right)}.$ (13)
@@
$k=\max\left\{1,\left\lceil\frac{16\sigma^{2}}{\Delta^{2}}\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\right\rceil\right\}$ (14)
@@
$R_{T}^{*}\leq\min\left\{T\Delta,\Delta+\frac{16\sigma^{2}}{\Delta}\left(1+\max\left\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\right\}\right)\right\}$ (15)
where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (13) with the value of the exploration time in (14). Then, focusing on $\max\left\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\right\}$, we obtain $\Delta\geq\frac{4\sigma}{\sqrt{T}}$. Thus, when we put it in the regret bound (15), we have the following:
@@
where $c>0$ is some constant. If $\Delta\leq 1$, it becomes $R_{T}^{*}\leq 1+c\sqrt{T}$, which refers to an instance-independent bound since it depends only on the time horizon $T$ not on the game instance. Our analysis shows that we obtain a regret bound comparable to the standard bandit setting in [20], indicating that the ETC-based algorithm performs effectively even in the zero-sum game setting.